OpenStudy (anonymous):

1. Solve: -6x - 3y = 30-3x + 6y = 0 a.(-4,-1) b.(-3,-1) c.(-4,-2) d,(-3,-2)

OpenStudy (anonymous):

OOOH

OpenStudy (anonymous):

just a sec, I keep coming up with wrong answers

OpenStudy (anonymous):

I keep getting x=-10. Really frustrating.

OpenStudy (anonymous):

-3y=30-3x+6x+6y -3y=30+3x+6y -6y-3y=30+3x -9y=30+3x y=-10/3 -1/3 -10/3 -1x/3 = 0 -1x/3 = 10/3 x=-10 ARRGH